# A Multi-Task Foundation Model for Wireless Channel Representation Using Contrastive and Masked Autoencoder Learning

### v1

@@ Abstract
Current applications of self-supervised learning to wireless channel representation often borrow paradigms developed for text and image processing, without fully addressing the unique characteristics and constraints of wireless communications. Aiming to fill this gap, we first propose WiMAE (Wireless Masked Autoencoder), a transformer-based encoder-decoder foundation model pretrained on a realistic open-source multi-antenna wireless channel dataset. Building upon this foundation, we develop ContraWiMAE, which enhances WiMAE by incorporating a contrastive learning objective alongside the reconstruction task in a unified multi-task framework. By warm-starting from pretrained WiMAE weights and generating positive pairs via noise injection, the contrastive component enables the model to capture both structural and discriminative features, enhancing representation quality beyond what reconstruction alone can achieve. Through extensive evaluation on unseen scenarios, we demonstrate the effectiveness of both approaches across multiple downstream tasks, with ContraWiMAE showing further improvements in linear separability and adaptability in diverse wireless environments. Comparative evaluations against a state-of-the-art wireless channel foundation model confirm the superior performance and data efficiency of our models, highlighting their potential as powerful baselines for future research in self-supervised wireless channel representation learning.
Index Terms:
Wireless Channel Modeling, Foundation Models, Transformer Architectures, Masked Autoencoder, Deep Learning.
@@ I Introduction
Large-scale self-supervised pretraining has transformed the fields of natural language processing and computer vision. This paradigm leverages diverse datasets and proxy objectives to learn broadly transferable representations, in contrast to traditional task-specific training approaches [2, 3, 4]. By decoupling feature learning from downstream tasks, it enables efficient, task-agnostic adaptation. Models following this two-stage strategy—computationally intensive pretraining followed by lightweight fine-tuning—are commonly referred to as foundation models [5].
@@ I-A Background and Motivation
Deep learning has demonstrated significant potential in wireless communications across a range of task-specific applications. For instance, deep neural networks have been employed for channel estimation and channel state information (CSI) feedback compression [6, 7, 8, 9, 10, 11], end-to-end communication system design [12, 13, 14], and generative channel modeling [15, 16, 17]. Similarly, classification-based tasks such as beam management [18, 19, 20] and link blockage prediction [21, 22] can benefit from supervised deep learning techniques. These advances are also reflected in ongoing standardization efforts within 3GPP [23].
However, the prevailing paradigm in the wireless domain remains task-specific, often requiring ad-hoc architectures, labels, and retraining for each use case. In contrast, the notion of foundation models—pretrained on large-scale data using self-supervised objectives and adaptable across tasks—has yet to be fully embraced in the wireless domain. A limited but growing body of work has begun to explore task-agnostic self-supervised learning for wireless channel representation, with most efforts falling into two broad methodological categories: contrastive and reconstructive learning.
@@
Contrastive learning typically utilizes encoder-only architectures to maximize representational similarity between augmented versions of identical data points (positive pairs) while minimizing similarity between distinct samples (negative pairs) [24, 25]. For example, one study treats CSI and channel impulse response (CIR) of the same wireless channel as positive pairs, maximizing their representational similarity in the embedding space [24]. A related investigation uses geospatial information by constructing a contrastive objective in which channels from identical geolocations serve as positive samples, while those from different locations serve as negative samples [25].
@@
In terms of encoder-decoder approaches, the work in [29] employs a vision transformer-based autoencoder for masked spectrogram prediction, demonstrating its effectiveness in human activity sensing and spectrogram segmentation tasks. Complementary research explores masked autoencoder pretraining for channel prediction [30], although it does not address transferability across different downstream tasks.
@@
Despite these advancements, key limitations remain in current wireless foundation models. BERT-based architectures such as LWM [33] typically employ low masking ratios (e.g., 15%), which reduce the complexity of the pretraining task and may result in shallow or overspecialized feature representation. Furthermore, several approaches adapt BERT-style models by replacing the original cross-entropy loss with regression objectives tailored to continuous-valued wireless data, without thoroughly reassessing the implications for representation learning. These design choices can constrain the expressiveness and generalizability of the learned representations, ultimately limiting performance across diverse downstream wireless tasks.
More broadly, while contrastive and reconstructive approaches have each shown promise, they face fundamental limitations when applied in isolation to wireless channel modeling. Contrastive methods, for instance, are sensitive to the choice of positive and negative samples—poor sample selection can mislead the model, especially in wireless scenarios where channels may share subtle but significant similarities. These methods also typically require large batch sizes to provide a sufficient number of negative samples, increasing computational overhead [34, 35]. In contrast, reconstructive approaches tend to emphasize the accurate recovery of low-level statistical patterns, which may come at the expense of not learning the discriminative features necessary for downstream tasks such as beam selection. This produces representations that work well for regression tasks but may underperform in classification or decision-making scenarios that depend on fine-grained feature distinctions [36]. Combining both approaches in a multi-task framework addresses these weaknesses by enabling simultaneous learning of both structural and discriminative channel features.
@@ I-B Contribution and Summary of Results
In this paper, we introduce Wireless Masked Autoencoder (WiMAE), a pretrained transformer-based wireless channel foundation model based on the MAE framework. WiMAE directly addresses the limitations of prior reconstructive approaches by adopting an asymmetric encoder-decoder architecture that emphasizes abstraction over memorization. Inspired by techniques in the vision domain [37, 26], WiMAE replaces the BERT-style regression loss with a structure that more effectively encourages the extraction of semantically meaningful features. Specifically, WiMAE’s encoder processes only the visible patches of the input, while a lightweight decoder reconstructs the masked portions. Unlike LWM’s shallow masking, WiMAE employs higher masking ratios, forcing the model to learn robust representations from limited observations and resulting in more transferable features across various wireless tasks.
Building upon the WiMAE foundation, we further enhance our approach by developing ContraWiMAE, which incorporates a contrastive learning objective alongside the reconstruction task in a unified multi-task framework. By warm-starting from pretrained WiMAE weights and generating positive pairs via the injection of additive white Gaussian noise (AWGN), ContraWiMAE contrasts these against negative samples drawn from different channel realizations. This enables the model to learn representations that are not only effective for reconstruction but also highly discriminative. The resulting hybrid approach significantly increases the linear separability of the learned embeddings, allowing simpler downstream models to achieve better performance without relying on complex architectures to compensate for representational weaknesses.
We evaluate the effectiveness of WiMAE and ContraWiMAE on two downstream tasks: cross-frequency beam selection and line-of-sight (LoS) detection. The beam selection task evaluates how well each encoding captures the channel’s detailed structure across antenna elements and subcarriers, while LoS detection measures the effectiveness of the extracted features as compact channel summaries. Key findings from our experiments include the following:
  • WiMAE achieves optimal performance at a masking ratio of 0.6, outperforming competing baselines by up to 36.5% increase in accuracy in linear probing tasks.
  • Increasing encoder depth yields diminishing returns beyond 12 layers, whereas deeper decoders improve representation quality without affecting inference latency.
  • Higher-dimensional embeddings improve performance, especially for complex tasks; increasing the representation dimension from 32 to 64 leads to substantial gains with diminishing returns beyond 64.
  • WiMAE demonstrates exceptional data efficiency, matching the performance of baseline models trained on complete datasets using only 1% of the training data.
  • While complex downstream models help reduce performance gaps for other approaches, WiMAE shows minimal improvement from simple to complex models—indicating that it already produces near-optimal representations.
  • The addition of a contrastive component in ContraWiMAE further enhances linear separability of the learned representations, leading to improved downstream performance, achieving up to 16.1% and 42.3% higher accuracy compared to WiMAE and other baselines, respectively.
  • Warm-starting from pretrained WiMAE enables efficient training of ContraWiMAE without degrading the strong reconstruction capabilities of the base model.
Relation to our prior work
This paper extends our preliminary work submitted to IEEE GLOBECOM 2025 [1]. While the conference version introduces the WiMAE architecture and demonstrates its promise for wireless channel representation learning via masked reconstruction, the present article offers substantial enhancements. Specifically: (i) we introduce ContraWiMAE, a novel multi-task extension that integrates contrastive learning to improve discriminability; (ii) we conduct a broader set of ablation studies and evaluate performance on two downstream tasks—beam selection and LoS detection—across a more exhaustive set of configurations, demonstrating significantly improved generalization, data efficiency, and linear separability of learned features; and (iii) we extract deeper insights into optimal architectural design, pretraining strategies, and fundamental relationships between model capacity, task complexity, and data efficiency.
@@
The remainder of this paper is organized as follows. Section II describes the system model and problem formulation, together with the training objectives used for model optimization. Section III presents the WiMAE and ContraWiMAE architectures, detailing the preprocessing steps, encoder and decoder designs, and the integration of contrastive learning. Section IV outlines the methodology, including dataset generation, pretraining setup, and adaptations for downstream tasks. Section V reports and thoroughly analyzes the experimental results, highlighting their implications for wireless representation learning. Finally, Section VI concludes the paper.
@@ II-A System Model
We consider a MISO-OFDM system with $N_{\text{f}}=32$ subcarriers, $30$ kHz subcarrier spacing, and a total system bandwidth of $B=960$ kHz. The base station utilizes a uniform linear array (ULA) with $N_{\text{s}}=32$ antennas and half-wavelength antenna spacing. The user equipment employs a single omnidirectional antenna. All antennas have isotropic radiation patterns and single polarization. The maximum number of propagation paths taken into account is set to 20 and Doppler effects are not considered. The frequency domain channel response for each symbol period is denoted as $\mathbf{H}\in\mathbb{C}^{N_{\text{s}}\times N_{\text{f}}}$ and is calculated between the base station and every user in a given scenario. To perform a fair comparison, our system model parameters are purposely chosen to match the configuration used in LWM [33].
@@ II-B Problem Formulation
We begin by formulating the basic WiMAE approach, then extend it to the ContraWiMAE framework that incorporates contrastive learning.
II-B1 WiMAE: Masked Reconstruction Objective
We model the problem with a parameterized encoder-decoder pair $(f_{\theta},g_{\phi})$. Given a complex channel matrix $\mathbf{H}\in\mathbb{C}^{N_{\text{s}}\times N_{\text{f}}}$, our aim is to learn an efficient representation that captures the underlying structure with only partial observations.
@@ II-B2 ContraWiMAE: Hybrid Reconstruction and Contrastive Learning
Building upon the WiMAE foundation, we enhance our approach by incorporating a contrastive learning objective alongside the reconstruction task in a multi-task framework.
@@
Let $\mathbf{H}_{i,\text{m}}=\mathbf{M}^{\prime}(\mathbf{H}_{i})$ and $\mathbf{H}_{i,\text{m}}^{+}=\mathbf{M}^{\prime}(\mathbf{H}_{i}^{+})$ represent the masked versions of the original and augmented channel matrices, respectively. Let $\mathbf{z}_{i}=\mathcal{G}_{\psi}(f_{\theta}(\mathbf{H}_{i,\text{m}}))$ and $\mathbf{z}_{i}^{+}=\mathcal{G}_{\psi}(f_{\theta}(\mathbf{H}_{i,\text{m}}^{+}))$ be the global embeddings derived by applying a nonlinear operator $\mathcal{G}_{\psi}$ parameterized by $\psi$, which generates a low-dimensional representation of the encoder output. These embeddings are normalized to lie on a unit hypersphere for contrastive loss calculation. Inspired by Noise Contrastive Estimation [38] and following the InfoNCE approach [39], we formulate the contrastive objective as:
$\begin{split}\mathcal{L}_{\text{contra}}(\theta,\psi)&=-\mathbb{E}_{\mathbf{H}_{i}\sim p(\mathbf{H}),\mathbf{M}^{\prime}\sim p(\mathbf{M}^{\prime})}\\ &\Big{[}\log\frac{\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{i}^{+}/\tau)}{\sum_{j=1,j\neq i}^{\mathcal{B}}\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{j}/\tau)}\Big{]},\end{split}$ (3)
where $\tau$ is a temperature parameter and $\mathcal{B}$ is the batch size. This loss encourages the representation of a channel and its augmented version to be similar, while pushing apart representations of different channels in the same batch.
@@ II-B3 Model Utilization Strategy
While our primary focus is on the encoder $f_{\theta}$, which produces the latent representations used for downstream tasks, both the decoder $g_{\phi}$ and the contrastive learning component $\mathcal{G}_{\psi}$ play critical roles during training. The decoder enforces the reconstruction objective, encouraging the encoder to extract structurally informative features. The contrastive learning objective enhances the discriminative properties of these representations by pulling similar channel realizations closer in the embedding space while pushing dissimilar ones apart.
This multi-task training approach allows ContraWiMAE to simultaneously learn both structural and discriminative features, leading to more robust representations that generalize better across diverse wireless environments. During inference, the decoder is typically discarded when only the encoder’s embeddings are needed (e.g., for downstream classification tasks). However, for tasks that require accurate signal reconstruction, the decoder remains relevant and can be retained to generate full outputs from partial observations.

### v2

@@ Abstract
Current applications of self-supervised learning to wireless channel representation often borrow paradigms developed for text and image processing, without fully addressing the unique characteristics and constraints of wireless communications. To bridge this gap, we introduce ContraWiMAE, Wireless Contrastive Masked Autoencoder, a transformer-based foundation model that unifies masked reconstruction and masked contrastive learning for wireless channel representation. Our key innovation is a new wireless-inspired contrastive objective that exploits the inherent characteristics of wireless environment, including noise, fading, and partial observability, as natural augmentation. Through extensive evaluation on unseen scenarios and conditions, we demonstrate our method’s effectiveness in multiple downstream tasks, including cross-frequency beam selection, line-of-sight detection, and channel estimation. ContraWiMAE exhibits superior linear separability and adaptability in diverse wireless environments, demonstrating exceptional data efficiency and competitive performance compared with supervised baselines under challenging conditions. Comparative evaluations against a state-of-the-art wireless channel foundation model confirm the superior performance and data efficiency of our approach, highlighting its potential as a powerful baseline for future research in self-supervised wireless channel representation learning. To foster further work in this direction, we release the model weights and training pipeline for ContraWiMAE.
@@ I Introduction
Large-scale self-supervised pretraining has transformed the fields of natural language processing and computer vision. This paradigm leverages diverse datasets and proxy objectives to learn broadly transferable representations, in contrast to traditional task-specific training approaches [2, 3, 4]. By decoupling feature learning from downstream tasks, it enables efficient, task-specific adaptation. Models following this two-stage strategy—computationally intensive pretraining followed by lightweight adaptation—are commonly referred to as foundation models [5].
@@ I-A Background and Motivation
Traditional task-specific approaches typically require extensive labeled datasets, substantial computational resources, and dedicated model storage for each application. These constraints become particularly limiting when task-specific data are scarce or when training computational resources are restricted. Moreover, AI/ML use cases are often developed in isolation, leading to high development costs and redundant data collection efforts where overlapping information is gathered separately, causing trained models to capture similar patterns independently.
Foundation models address these challenges through task-agnostic learning that employs a single backbone model to provide shared representations, with dedicated heads designed for specific use cases. This framework enables efficient adaptation to diverse downstream applications with minimal additional data and computation, as existing knowledge captured in the shared backbone can be reused across tasks [6]. The approach simplifies monitoring and management since maintaining one unified model is considerably easier than handling multiple specialized ones, and retraining or fine tuning can be accomplished within a single pipeline.
This paradigm proves especially valuable in scenarios such as new frequency band deployments where historical data are limited, edge computing environments with constrained computational resources, and dynamic spectrum access systems requiring rapid adaptation to varying channel conditions. Furthermore, when complementary tasks on the user equipment side, such as beam management or channel state information (CSI) compression, share portions of the backbone architecture, task-agnostic models facilitate the integration of two-sided training or inference routines, thereby supporting enhanced model interoperability and system-wide optimization.
@@
Despite these compelling advantages, the prevailing paradigm in the wireless domain remains task-specific, with each application requiring dedicated architectures and training. The task-specific approach has shown significant potential in applications such as channel estimation and CSI feedback compression [7, 8, 9, 10, 11, 12], end-to-end communication system design [13, 14, 15], and generative channel modeling [16, 17, 18]. Similarly, classification-based tasks such as beam management [19, 20, 21] and link blockage prediction [22, 23] benefit from supervised deep learning techniques. These advances are also reflected in ongoing standardization efforts within 3GPP [24]. However, while these task-specific successes validate the potential of AI/ML in wireless systems, they also highlight the opportunity for more unified approaches that can leverage shared channel structure across multiple applications.
We argue that the wireless communication domain presents unique opportunities for the foundation model development, as CSI inherently contains rich structural information that can benefit numerous downstream tasks. Specifically, CSI matrices serve as the fundamental information substrate that drives most critical decisions in modern communication systems, from beamforming and precoding to resource allocation and interference management. This convergence on a single input modality makes the wireless domain particularly well suited for foundation models, as learning representations directly from CSI enables knowledge transfer across diverse tasks that traditionally require separate specialized models.
To capitalize on this opportunity, we propose ContraWiMAE  [^1] We make ContraWiMAE along with our dataset publicly available at: https://github.com/BerkIGuler/WirelessContrastiveMaskedLearning (Wireless Masked Contrastive Autoencoder), a multi-task learning framework that introduces a novel masked contrastive objective into a wireless masked autoencoder foundation. Our approach leverages the inherent complexity of wireless channels, including noise, fading, and partial observability, as natural augmentation, contrasting differently masked versions of the same channel to learn representations that capture both structural coherence and discriminative features. In parallel with our efforts, a limited but growing body of work has begun to explore task-agnostic self-supervised learning for wireless channel representation, with most efforts falling into two broad methodological categories: contrastive and reconstructive learning.
@@
Contrastive learning typically utilizes encoder-only architectures to maximize representational similarity between augmented versions of identical data points (positive pairs) while minimizing similarity between distinct samples (negative pairs) [25, 26]. For example, one study treats the CSI and the channel impulse response (CIR) of the same wireless channel as positive pairs, maximizing their representational similarity in the embedding space [25]. A related investigation uses geospatial information by constructing a contrastive objective in which channels from identical geolocations serve as positive samples, while those from different locations serve as negative samples [26].
@@
In terms of the encoder-decoder approaches, the work in [30] employs a vision transformer-based autoencoder for masked spectrogram prediction, demonstrating its effectiveness in human activity sensing and spectrogram segmentation tasks. Complementary research explores masked autoencoder pretraining for channel prediction [31], although it does not address transferability across different downstream tasks.
@@
Despite these advancements, key limitations remain in current wireless foundation models. BERT-based architectures such as LWM [34] typically employ low masking ratios (e.g., 15%), which reduce the complexity of the pretraining task and may result in shallow or overspecialized feature representation. Furthermore, several approaches adapt BERT-style or GPT-style models by replacing the original cross-entropy loss, designed for prediction over a discrete set, with regression objectives tailored to continuous-valued wireless data, without thoroughly reassessing the implications for representation learning. These design choices can limit the expressiveness and generalizability of the learned representations, ultimately limiting the performance of diverse downstream wireless tasks.
More broadly, while contrastive and reconstructive approaches have each shown promise, they face fundamental limitations when applied in isolation to wireless channel modeling. Contrastive methods, for instance, are sensitive to the choice of positive and negative samples—poor sample selection can mislead the model, especially in wireless scenarios where channels may share subtle but significant similarities. These methods also typically require large batch sizes to provide a sufficient number of negative samples, increasing computational overhead [35, 36]. In contrast, reconstructive approaches tend to emphasize the accurate recovery of low-level statistical patterns, which may come at the expense of not learning the discriminative features necessary for downstream tasks such as beam selection. This produces representations that work well for regression tasks, but may underperform in classification or decision-making scenarios that depend on fine-grained feature distinctions [37]. Combining both approaches in a multi-task framework addresses these weaknesses by enabling simultaneous learning of both structural and discriminative channel features.
@@ I-B Contribution and Summary of Results
Our approach directly addresses the limitations of prior reconstructive approaches by adopting an asymmetric encoder-decoder architecture that emphasizes abstraction over memorization. Inspired by techniques in the vision domain [38, 27], we utilize the regression loss with a structure that more effectively encourages the extraction of semantically meaningful features compared to encoder-only and decoder-only architectures that naively substitute regression for classification losses. Specifically in our approach, the encoder only processes the visible patches of the input, while a lightweight decoder reconstructs the masked portions from the latent representations of visible patches.
This reconstruction-only foundation serves as both a strong baseline for comparison and a critical building block for our complete framework. Building upon this reconstruction-only foundation, we introduce a novel masked contrastive objective and continually pretrain with a multi-task objective. ContraWiMAE, instead of requiring engineered augmentations or heuristic-based positive sample selection, generates contrastive pairs by applying different random masks to noisy versions of identical channel realizations. The key innovation lies in teaching the model to recognize that different partial views of the same propagation environment should map to similar representations, while maintaining discriminative boundaries between fundamentally different channels. This masked contrastive objective enables the extraction of invariant features that remain consistent across varying noise conditions and measurement patterns, which are critical properties for real wireless deployments where only noisy and partial channel observations are available. Furthermore, ContraWiMAE preserves the strong reconstruction capabilities through multi-task learning while significantly enhancing the linear separability of learned embeddings.
We evaluate our approach on three complementary downstream tasks that represent different categories of wireless applications: cross-frequency beam selection (a complex classification task requiring detailed spatial understanding), line-of-sight detection (a simple classification task for coarse channel characterization), and channel estimation (a reconstruction task leveraging learned channel structure). Although our current evaluation focuses on these representative tasks, the task-agnostic nature of our learned representations provides a foundation for broader applicability to additional wireless communication challenges.
Key findings from our experiments include the following:
  • Our dual masked contrastive and reconstructive learning objective enables competitive performance with supervised training across tasks.
  • Masked contrastive learning using noise as augmentation improves robustness to low-SNR conditions and distribution shifts across frequencies and environments.
  • The masked contrastive objective promotes a linearly separable and locally structured feature space, demonstrated by linear probing and nearest-neighbor evaluations.
  • ContraWiMAE enables simple downstream models to match the performance of complex networks, enabling practical wireless adoption.
  • Our dual learning strategy maintains reconstruction capabilities with minimal degradation compared to reconstruction-only learning.
@@
The remainder of this paper is organized as follows. Section II describes the system model and problem formulation, together with the training objectives used for model optimization. Section III presents the ContraWiMAE architecture, detailing the preprocessing steps, encoder and decoder designs, and the integration of contrastive learning. Section IV outlines the methodology, including dataset generation, pretraining setup, and adaptations for downstream tasks. Section V includes the analysis of the computational complexity of our method. Section VI reports and thoroughly analyzes the experimental results, highlighting their implications for wireless representation learning. Finally, Section VII concludes the paper.
@@ II-A System Model
We consider a MISO-OFDM system with $N_{\text{f}}$ subcarriers and $\Delta f$ kHz spacing where the base station utilizes a uniform linear array (ULA) with $N_{\text{s}}$ antennas. The user equipment employs a single omnidirectional antenna. The channel response in the frequency domain for each symbol period is denoted by $\mathbf{H}\in\mathbb{C}^{N_{\text{s}}\times N_{\text{f}}}$ and represents the channel between the base station and the users. The assumption of available channel state information follows standard practice in machine learning for wireless communications research, where CSI is typically obtained through pilot-based channel estimation in contemporary systems like LTE and 5G. To perform a fair comparison, our system model parameters are purposely chosen to match the configuration used in LWM [34].
@@ II-B Problem Formulation
We begin by formulating the reconstruction approach, then extend it to the ContraWiMAE framework that incorporates our masked contrastive objective. Our unconstrained pretraining objectives are designed to learn generic representations without bias toward specific constraint sets, enabling task-specific constraints to be incorporated during downstream application phases.
II-B1 Masked Reconstruction Objective
We model the problem with a parameterized encoder-decoder pair $(f_{\theta},g_{\phi})$. Given a complex channel matrix $\mathbf{H}\in\mathbb{C}^{N_{\text{s}}\times N_{\text{f}}}$, our aim is to learn an efficient representation that captures the underlying structure with only partial observations. We employ random masking rather than structured patterns to learn general-purpose representations that adapt to various loss patterns without bias toward specific failure modes. This promotes robust feature learning that transfers to actual structured losses during downstream applications.
@@ II-B2 ContraWiMAE: Hybrid Reconstruction and Contrastive Learning
We enhance our approach by incorporating a contrastive learning objective alongside the reconstruction task in a multi-task framework.
@@
[…]
$\begin{split}\mathcal{L}_{\text{contra}}(\theta,\psi)&={\color[rgb]{0,0,0}-\frac{1}{2\mathcal{B}}\sum_{k=1}^{2\mathcal{B}}\log\frac{\exp(\mathbf{z}_{k}\cdot\mathbf{z}_{\text{pos}(k)}/\tau)}{\sum_{j=1,j\neq k}^{2\mathcal{B}}\exp(\mathbf{z}_{k}\cdot\mathbf{z}_{j}/\tau)}},\end{split}$ (3)
where $\mathbf{z}_{k}$ denotes the $k$-th row of $\mathbf{Z}$, $\text{pos}(k)=k+\mathcal{B}$ if $k\leq\mathcal{B}$, otherwise $\text{pos}(k)=k$-$\mathcal{B}$, $\tau$ is a temperature parameter, and $\mathcal{B}$ is the batch size.
@@ II-B3 Model Utilization Strategy
While our primary focus is on the encoder $f_{\theta}$, which produces the latent representations used for downstream tasks, both the decoder $g_{\phi}$ and the contrastive learning component $\mathcal{G}_{\psi}$ play critical roles during training. The decoder enforces the reconstruction objective, encouraging the encoder to extract structurally informative features. The masked contrastive objective favors representations that are invariant to particular antennas and subcarriers while remaining noise invariant. This prevents overfitting to noise artifacts and specific measurement patterns, while maintaining semantic consistency across partial observations.
This multi-task training approach allows ContraWiMAE to simultaneously learn both structural and discriminative features, leading to more robust representations that generalize better across diverse wireless environments. During inference, the decoder and contrastive head are typically discarded when only the encoder’s embeddings are needed (e.g., for downstream classification tasks). Therefore, the decoder complexity may not pose as serious implications as encoder complexity for inference time and compute resources. However, for tasks that require accurate signal reconstruction, the decoder remains relevant and can be retained to generate full outputs from partial observations.